OpenStudy (anonymous):

The radius of the Earth is apprximately 3,960 miles, while the radius of the Moon is approximately 1,080 miles. How many times larger is the surface area of the Earth than the surface area of the Moon? A. 3.7 B. 11.1 C. 13.4 D. 50

OpenStudy (johnweldon1993):

Well we can think of both the objects as spheres... the surface area of a sphere is \(\large 4 \pi r^2\) So lets see...the radius of the earth....is \(\large \frac{3960}{1080} \approx 3.7\) times bigger than the moons radius So say the moons radius was 1 ...we would have \(\large 4 \pi\) And the earths would then be \(\large 4 \pi (3.7)^2 = 4 (13.44)\pi\) So it looks like the surface area of the earth is 13.4 times bigger than the surface area of the moon

OpenStudy (anonymous):

ooh that is what i did wrong i didn't square the 4 ha ha ha. Thanks for the help i got it!